OpenStudy (anonymous):

Suppose that a and b vary inversely and that b=5/4 when a=8. Write a function that models the inverse variation and find b when a =15.

OpenStudy (anonymous):

@mukushla @mathmale @e.mccormick

OpenStudy (anonymous):

well a vary inversely with variations of b, that means you can write:\[a =\frac{k}{b}\]is it clear?

OpenStudy (anonymous):

Yes, it is.

OpenStudy (anonymous):

can you find constant \(k\)

OpenStudy (anonymous):

Don't you do a*b=k?

OpenStudy (anonymous):

use the fact that b=5/4 when a=8 :-) and find \(k\)

OpenStudy (anonymous):

Ok thank you! So would it be 10? would K equal 10?

OpenStudy (anonymous):

Oh wait, did I do this right?

OpenStudy (anonymous):

@mukushla ?

OpenStudy (anonymous):

@ganeshie8 @Hero ?

hero (hero):

ab = k 8(5/4) = k 10 = k Yes, 10 is correct so far.

OpenStudy (anonymous):

Ok then do I plug in 15 for a?

hero (hero):

Basically, from here you use formula ab = k again This time, k = 10 (still), but a = 15. Input those values then solve for b

OpenStudy (anonymous):

Okay thank you, this clears it all up!

hero (hero):

You're welcome

OpenStudy (anonymous):

So I got b=5/4a

OpenStudy (anonymous):

This is when I simplified it, it's just that its asking me to write a function.

hero (hero):

You input 15(b) = 10 and ended up with b = 5/4a? How?

OpenStudy (anonymous):

It's looking for the function.

hero (hero):

At some point though, you got confused.

OpenStudy (anonymous):

Oh... that makes sense.

hero (hero):

Basically you're suggesting that b = 5/4a is the same as ab = 5/4 You're suggesting that 5/4 = k which isn't true.

hero (hero):

The only consistent value is k = 10. You want to include that in your formula. But one thing I notice is, you never solved for b

OpenStudy (anonymous):

I didn't?

hero (hero):

No. You were supposed to take ab = k (the formula for inverse variation) then input a = 15 and k = 10 15b = 10 Then solve for b Just try doing that please.

OpenStudy (anonymous):

So that would mean it's b=10/a! @Hero

hero (hero):

Yes, but you still haven't solved for b. That's the formula though

OpenStudy (anonymous):

Thank you :) I got .666

hero (hero):

b needs to be in fraction form

OpenStudy (anonymous):

10/15

OpenStudy (anonymous):

b=10/15

hero (hero):

In simplified fraction form

OpenStudy (anonymous):

Oh! 1/3

OpenStudy (anonymous):

Wups 2/3

hero (hero):

How did you get from 10/15 to 1/3? Show your work

hero (hero):

Yes, correct b = 2/3 and the formula is ab = 10

OpenStudy (anonymous):

Simplified by 5, okay thanks for clearing this up @Hero

hero (hero):

For some reason this particular topic confuses 90% of students

hero (hero):

It's not even that hard

OpenStudy (anonymous):

But for some reason I get a little confused sometimes

OpenStudy (anonymous):

Thanks again :)

hero (hero):

yw